Sidebar: Functional PageRank (PR)

Computing PageRank (PR)

PageRank as a random surfer process: Start surfing from a random node and keep following links with probability μ restarting with probability 1 − μ; the node for restarting will be selected based on a personalization vector v. The ranking value x_{i} of a node i is the probability of visiting this node during surfing.

PR can also be cast in power series representation as x = (1 − μ) ∑_{j=0}^{k} μ^{j}S^{j}v; S encodes column-stochastic adjacencies.

Functional rankings

A general method to assign ranking values to graph nodes as x = ∑_{j=0}^{k} ζ_{j}S^{j}v. PR is a functional ranking, ζ_{j} = (1 − μ)μ^{j}.

Terms attenuated by outdegrees in S and damping coefficients ζ_{j}.

Ananth Grama (Purdue University)  TransMR: Data-Centric Programming  1 / 1